If the electric potential varies as ? = 3 ?^2 ? + 2 ? ?^2  , where x and y are both expressed in meters, determine:

a) an algebraic expression for the x and y components of the electric field

b) numerical values for the components of the electric field at the point (2, 5) ?
Answer #1

-:utos Electric field ( ² ) =-ov E-13 + 1) м - -{ x it i) ) he иеле у 2 (3x+2+2х) = kx + 2-3- ә ә) ə V с (3x+2+2xy-) = 3х-+ 4